# Smoothed Graph Contrastive Learning via Seamless Proximity Integration

### v1

@@ D.1 Node classification in full-batch scenario
SGCL-B 85.08±0.12 72.77±0.33 83.67±0.06 92.16±0.15 88.24±0.05 92.43±0.03
SGCL-D 84.47±0.25 70.32±0.04 85.22±0.02 92.04±0.05 84.98±0.34 90.09±0.11
While SGCL continues to demonstrate superior performance across most benchmarks in this case, it is noteworthy that SGCL exhibits enhanced performance in the mini-batch scenario compared to the full-batch. We attribute this difference to the presence of a substantial number of connected components in the full-batch graph. However, our framework benefits from mini-batch graphs, which can generate more connected mini-batches using a random-walk sampling approach. This enables our framework to leverage proximity information within the mini-batch graphs more effectively than in the full-batch graph.
D.2 Computational analysis
@@
The computational cost of graph contrastive learning models is analyzed through two distinct components: pre-training and downstream task evaluation. In the pre-training phase, the process involves augmentation generation, encoder computation, and computation of the contrastive loss function for each batch. In the downstream task phase, the model learns two input/output MLP layers and evaluates the model for node classification. We conduct the computation analysis to evaluate the runtime performance of three variants of the SGCL model and compare these variants with several baseline methods on graphs of different scales, ranging from small to medium and large-scale graphs. The results of these experiments are summarized in Table 7.
@@
[…]
@@
These results indicate that during pre-training, SGCL-T on the small-scale graph outperforms MVGRL in running time. For medium-scale and large-scale graphs, the computational are approximately 18% and 40% higher than those of MVGRL, respectively. The computational cost of the other variants of the model is increased compared to the baselines. However, the observed increase in computations aligns with the expectations, as SGCL integrates supplementary information into the conventional contrastive loss function, and the smoothing strategies require the exploration of the graph to identify proximity information.
It’s noteworthy to highlight that the computational costs in the downstream evaluation phase across all models are nearly identical on each benchmark. This implies that, despite the more computations during the pre-training phase, our model demonstrates efficiency during the downstream evaluation phase.
Appendix E Ablation study
E.1 Evaluating with other mini-batching generation methods
@@
SGCL-T RW-sampler 86.54±1.4 73.71±2.0 86.57±0.49 92.99±0.36 87.83±1.4 93.05±0.94
Ego-graph 84.93±4.2 71.20±3.4 85.70±0.84 93.09±0.37 86.38±0.9 93.15±0.86
Node-sampler 84.12±1.8 70.42±3.3 85.94±0.49 92.87±0.58 85.92±0.44 92.24±0.65
Edge-sampler 85.59±2.1 70.66±1.4 86.76±0.43 93.23±0.86 86.28±1.3 92.31±1.2
[…]
Ego-graph 84.56±1.3 72.87±1.7 85.88±0.33 93.10±0.44 86.57±1.2 93.05±0.85
Node-sampler 84.49±1.1 70.84±1.9 85.90±1.0 93.18±0.4 85.68±0.57 91.95±1.0
Edge-sampler 84.26±1.7 71.86±2.2 86.19±0.7 92.85±0.64 84.52±0.58 91.66±1.1
[…]
Ego-graph 85.00±1.4 72.57±2.4 85.75±0.98 92.93±0.3 80.04±0.82 89.07±0.71
Node-sampler 84.26±1.7 72.51±1.1 85.18±1.0 93.06±0.41 78.68±0.64 89.33±1.6
Edge-sampler 86.47±1.2 70.42±0.83 86.57±0.7 93.26±0.08 73.31±0.74 90.61±0.32
E.2 Influence of different terms in contrastive loss function
Since the number of non-zero values in $\tilde{\Pi}_{neg}^{(i,j)}$ exceeds those in $\tilde{\Pi}_{pos}^{(i,j)}$, we initially assign $\lambda$ as $1/2N$. This adjustment aims to achieve a trade-off between positive and negative pairs within the loss function 5. However, in the experiments, we determined its optimal value through grid search. For instance, on the Photo dataset, the optimal value for $\lambda$ was found to be around $42.3e-$ 4. This value aligns with our first initialization when considering the batch size of $N=2000$ in the experiments.
To perform an ablation study on the contrastive loss function, we evaluate the significance of each term of Equation 5 and subsequently combine them with hyperparameter $\lambda$. Table E.2 provides the accuracies of different variants of SGCL achieved by different components of the contrastive loss function on three benchmarks of varying scales: small (Cora), medium (CoauthorCS), and large (ogbn-arxiv). Initially, we observe that the exclusion of any term from the loss function results in deteriorated or collapsed solutions, aligning with our expectations. Subsequently, we investigated the influence of the combination of two individual terms using an optimal value of $\lambda$.
@@ E.2 Influence of different terms in contrastive loss function
Model Benchmark (A) (B) $\mathcal{L}_{SGCL}^{(i,j)}$ $(\lambda)$
small (Cora) 84.13±3.1 83.54±1.0 86.54±1.4 (4e-4)
SGCL-T medium (CoauthorCS) 91.36±0.7 90.87±0.4 92.99±0.4 (1e-4)
@@
[…]
SGCL-B medium (CoauthorCS) 93.04±0.1 91.45±0.3 93.15±0.4 (1e-4)
@@
[…]
SGCL-D medium (CoauthorCS) 92.4±0.52 91.09±0.2 93.2±0.4 (1e-4)
@@
[…]
(B): $\parallel(1-\tilde{\Pi}_{pos}^{(i,j)})\odot\mathbf{C}^{(i,j)}\parallel_{F}^{2}$

### v2

@@ E.2 More evaluation on heterophilic graphs
Graph #Nodes #Edges
#Features
#Classe Class types $h$
Chameleon 2,277 36,101 2,325 5 Wiki pages 0.23
Actor 7,600 29,926 931 5 Actors in movies 0.22
Cornell 183 295 1,703 5 Web pages 0.3
Texas 183 309 1,703 5 Web pages 0.11
Wisconsin 251 499 1,703 5 Web pages 0.21
Genius 421,961 984,979 12 2 marked act. 0.618
Twitch-gamers 168,114 6,797,557 7 2 mature content 0.545
We evlauate the performance of SGCL on these graphs and compare the results with conventional GCL approaches like the GRACE model. Results in Table 8 indicate that the proposed models still outperform conventional GCL on heterophilic graphs. However, the advantage of smoothing methods in homophilic graphs becomes more pronounced. As the homophily rate increases, the number of false negatives also increases, emphasizing the critical role of SGCL in effectively contributing both positive and negative pairs to the contrastive loss.
Table 8: Comparison of the accuracy of proposed SGCL models on heterophilic graphs with GCL.
Model Chameleon Actor Cornell Texas Wisconsin Genius Twitch-gamers
GCL (GRACE) 45.3±0.7 28.55±0.28 52.21±1.3 53.07±2.4 51.60±2.4 81.48±0.15 58.21±0.51
SGCL-T 46.32±1.3 29.29±0.47 54.45±0.6 54.56±1.8 52.50±2.3 82.58±0.33 59.37±0.47
SGCL-B 45.52±0.3 29.15±0.28 55.79±1.2 54.97±2.3 50.23±2.2 82.47±0.17 59.38±0.88
SGCL-D 45.91±2.5 28.74±0.76 53.12±5.4 55.48±2.7 53.62±2.1 81.86±0.2 58.60±0.48
E.3 Computational analysis
@@
The computational cost of graph contrastive learning models is analyzed through two distinct components: pre-training and downstream task evaluation. The pre-training phase consists of mini-batch generation, augmentation generation, encoder computation, and computations of the smoothing strategy. In the downstream task phase, the model learns two input/output MLP layers and evaluates the model for tasks such as node classification.
For a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with $N$ nodes and $E$ edges, the encoder computation using a message-passing-based GNN encoder $f_{\theta}$ efficiently computes embeddings with the complexity of $\mathcal{O}(N+E)$. The computation cost of graph augmentation consists of applying the feature mask ($\mathcal{O}(N)$) and the edge removal mask ($\mathcal{O}(E)$). The overall complexity is $\mathcal{O}(N+E)$, with the edge removal mask being the dominant factor ($\mathcal{O}(E)$). Notably, this cost is lower than that of MVGRL [12], as it utilizes a Personalized PageRank-based graph diffusion approach for structural augmentations, which entails a complexity of $\mathcal{O}(I.E)$, where $I$ represents the number of iterations required for convergence.
The proposed model imposes additional computational overhead compared to the standard GCL. This includes the computation associated with the mini-batch strategy and integrating the smoothing strategy into the conventional contrastive loss.
The computation of mini-batch strategies can be disregarded in the overall complexity analysis, as it is performed offline during preprocessing. However, the complexities of mini-batch strategies to generate $k$ subgraphs with a batch size $m$ from a given graph are $\mathcal{O}(km+kE)$ for the Random Node Sampler, $\mathcal{O}(km+kE)$ for the Random Edge Sampler, $\mathcal{O}(k(r\times h×d+m))$ for the Random Walk Sampler with $r$ roots, walk length $h$ and average node degree $d$, and $\mathcal{O}(k(s×d+m))$ for the $s$-hop Ego-graph sampler.
The main computational overhead is associated with the smoothing strategy. Taubin smoothing, which utilizes Laplacian matrices, has a computational complexity of $\mathcal{O}(N+E)$ constructing the Laplacian matrix and $\mathcal{O}(N)$ per iteration for matrix multiplication. The overall complexity, influenced by the number of iterations, ranges from linear to quadratic concerning the number of nodes $N$ and edges $E$.
The computational complexity of bilateral smoothing, which considers both spatial proximity and intensity similarity for each node, is predominantly influenced by the node degrees and the total number of edges, typically in the order $\mathcal{O}(N+E)$.
The computational complexity of diffusion-based smoothing primarily depends on the number of nodes $N$ and the number of iterations $K$. Each iteration involves summing the values of neighboring nodes, which can be considered $\mathcal{O}(deg(i))$ for each node $i$, where $deg(i)$ is the degree of node $i$. Therefore, the overall computational complexity can be expressed as $\mathcal{O}(K×N+E)$, where $K$ is the number of iterations, $N$ is the number of nodes, and $E$ is the total number of edges in the graph.
@@
GRLC pre-training 0.1193 0.3249 0.5747
downstream 0.0682 0.2685 0.4325
ProGCL-weight pre-training 0.0929 0.3428
downstream 0.0032 0.0152
ProGCL-mix pre-training 0.1192 0.4993
downstream 0.0029 0.0152
GREET pre-training 0.1793 2.5811
downstream 0.0031 0.0125
SGCL-T pre-training 0.1133 0.9723 1.3921
@@
For numerical evaluation, we conduct the computational analysis to evaluate the runtime performances of three variants of the SGCL model, comparing them with several baseline methods across graphs of varying scales. The results of these experiments are summarized in Table 9. For this analysis, we included the majority of baselines for which implemented code was accessible; benchmarks with unreasonable runtimes are marked with a dash (–) in the table.
These results indicate that during pre-training, SGCL-T on the small-scale graph outperforms MVGRL, GRLC, ProGCL-mix, and GREET in running time. For medium-scale and large-scale graphs, the computational costs are approximately 18% and 40% higher than those of MVGRL, respectively. The computational cost of the other variants of the model is increased compared to the baselines. This observed computational overhead is associated with the expectations, as SGCL integrates supplementary information into the conventional contrastive loss function, and the smoothing strategies require the exploration of the graph to identify proximity information.
Notably, the run-time performances in the downstream evaluation phase across baselines implemented in the PyGCL library, which follows a framework similar to SGCL, are nearly identical on each benchmark. This implies that, despite the more computation time in the pre-training phase, our model performs effectively in the downstream evaluation phase.
Additionally, we have conducted memory consumption comparisons for these baselines across graphs of different sizes, employing various graph batch generation methodds: Random Walk Sampler (RWS), Ego Graph Sampler (EGS), Random Node Sampler (RNS), and Random Edge Sampler (RES). The reported values reflect the GPU memory required (in MB) to train the models.
Table 10 indicate that the proposed SGCL models utilizing NSP and ESP demonstrate better memory efficiency, outperforming most baselines, particularly on medium and large-scale graphs. The EGS strategy tends to be more memory-intensive than other sampling methods. This becomes a crucial consideration when working with small and medium-scale graphs, particularly when comparing our approach to more memory-efficient baselines such as DGI, BGRL, and GBT.
Table 10: Memory consumption comparison between the proposed model and baselines on graphs of varying scales (MB).
Model
Small
Medium
Large
(ogbn-arxiv)
DGI 711 3023 15993
GRACE 1447 34255 34907
MVGRL 2753 33641 35191
BGRL 901 9931 26691
GBT 971 5093 26943
GRLC 1139 11533 35263
ProGCL-weight 1113 14129
ProGCL-mix 1647 28239
GREET 1655 33167
SGCL-T RWS 1957 11629 11639
EGS 2097 32782 38641
NSP 1075 2001 1489
ESP 1481 4249 3605
SGCL-B RWS 1973 9113 11293
EGS 2529 32137 38759
NSP 1075 1999 1473
ESP 1481 4321 4497
SGCL-D RWS 1973 11409 12051
EGS 2119 32871 35725
NSP 1977 2033 1477
ESP 1477 4715 3163
Appendix F Ablation study
F.1 Evaluating with other mini-batching generation methods
@@
SGCL-T RW-sampler 84.33±0.4 74.94±0.8 84.25±0.3 92.25±0.1 87.21±0.4 93.12±0.7
Ego-graph 84.21±0.3 73.88±1.6 84.47±0.7 92.12±0.3 86.7±0.6 93.05±0.7
Node-sampler 84.12±0.8 74.12±1.3 84.14±0.4 92.17±0.7 87.08±0.4 92.84±1.2
Edge-sampler 84.53±.5 73.54±1.7 83.76±0.6 91.83±0.8 86.88±1.5 93.33±0.8
SGCL-B RW-sampler 84.78±0.3 74.30±1.4 84.1±0.2 92.33±0.4 89.75±0.8 93.72±0.1
Ego-graph 84.63±0.7 73.26±0.7 84.16±0.5 91.68±0.7 89.07±1.6 93.13±0.4
Node-sampler 84.39±0.8 73.74±1.4 83.84±1.1 92.18±0.3 88.25±0.8 92.22±0.6
Edge-sampler 84.11±1.6 74.22±1.5 83.79±0.5 92.22±0.6 88.84±0.2 92.16±1.2
SGCL-D RW-sampler 84.17±0.4 75.72±0.8 85.12±0.3 92.14±0.2 86.11±0.3 92.87±0.6
Ego-graph 84.15±1.5 73.87±0.6 84.73±0.2 92.17±0.6 84.24±0.5 92.26±0.5
Node-sampler 84.23±1.3 74.43±1.3 84.68±1.3 92.05±0.3 85.38±0.4 91.63±1.4
Edge-sampler 84.75±1.4 73.55±0.8 85.17±0.6 92.21±0.1 85.3±0.7 93.76±0.2
F.2 Influence of different terms in contrastive loss function
Since the number of non-zero values in $\tilde{\Pi}_{\text{neg}}^{(i,j)}$ exceeds those in $\tilde{\Pi}_{\text{pos}}^{(i,j)}$, we initially assign $\lambda$ as $1/2N$. This adjustment aims to achieve a trade-off between positive and negative pairs within the loss function 5. However, in the experiments, we determined its optimal value through grid search. For instance, on the Photo dataset, the optimal value for $\lambda$ was found to be around $42.3e-$ 4. This value aligns with our first initialization when considering the batch size of $N=2000$ in the experiments.
To perform an ablation study on the contrastive loss function, we evaluate the significance of each term of Equation 5 and subsequently combine them with hyperparameter $\lambda$. Table 12 provides the accuracies of different variants of SGCL achieved by different components of the contrastive loss function on three benchmarks of varying scales: small (Cora), medium (CoauthorCS), and large (ogbn-arxiv). Initially, we observe that the exclusion of any term from the loss function results in deteriorated or collapsed solutions, aligning with our expectations. Subsequently, we investigated the influence of the combination of two individual terms using an optimal value of $\lambda$.
@@ F.2 Influence of different terms in contrastive loss function
Model Benchmark (A) (B) $\mathcal{L}_{\text{SGCL}}^{(i,j)}$ $(\lambda)$
small (Cora) 84.12±0.7 83.24±1.2 84.33±0.4 (4e-4)
SGCL-T medium (CoauthorCS) 92.15±0.3 91.74±0.4 92.25±0.1 (1e-4)
@@
small (Cora) 84.27±0.8 83.84±0.7 84.78±0.3 (4e-4)
SGCL-B medium (CoauthorCS) 91.73±0.4 91.66±0.7 92.33±0.4 (1e-4)
@@
small (Cora) 83.9±1.3 83.82±1.4 84.17±0.4 (4e-4)
SGCL-D medium (CoauthorCS) 91.63±0.4 91.32±0.6 92.14±0.2 (1e-4)
@@
(A): $\parallel{\tilde{\Pi}_{\text{pos}}^{(i,j)}\odot(1-\mathbf{C}^{(i,j)})}\parallel_{F}^{2}$
(B): $\parallel(1-\tilde{\Pi}_{\text{pos}}^{(i,j)})\odot\mathbf{C}^{(i,j)}\parallel_{F}^{2}$
F.3 Ablation analysis of hyperparameters in smoothing approaches
F.3.1 Taubin smoothing
Taubin smoothing is a combined process that alternates between negative and positive Laplacian filters to smooth the signal. The negative Laplacian filter, with hyperparameter $\mu<0$, smooths the input matrix $\mathbf{V}$, while the positive Laplacian filter, with hyperparameter $\tau>0$ (where $\mu<-\tau)$, prevents oversmoothing by restoring some of the original values. The smoothing process alternates $K$ times between these two filters. The values of $\tau$, $\mu$, and $K$ are carefully chosen to balance the smoothing process with the positive correction, ensuring a stable result that avoids both excessive noise and oversmoothing.
For the ablation study, we begin by fixing $K=2$ and evaluate the impact of varying the hyperparameters $\tau$ (where $10<\tau<$ 1) and $\mu$ (with $\mu<-\tau$). For a more detailed evaluation, we focus on the interaction between the parameters $\tau$ and $\mu$, both of which are varied over finer increments. Specifically, we evaluate the performance of the model on the Cora and Pubmed by adjusting $\tau$ in the range of $0.1$ to $0.5$ and $\mu$ between $-0.2$ and $-0.6$ $(\mu<-\tau)$. Table 13 highlights the results for different combinations of $\tau$ and $\mu$.
Table 13: Performance of SGCL-T with fixed $K=2$ and different combinations of $\mu$ and $\tau$.
Cora Pubmed
$\mu$0.2 -0.3 -0.4 -0.5 -0.6 -0.2 -0.3 -0.4 -0.5 -0.6
$\tau$ 0.1 84.87 84.31 84.06 84.21 83.71 83.9 83.22 83.54 82.56 81.82
0.2 - 83.55 84.31 84.37 84.03 - 83.36 84.32 82.22 82.06
0.3 - - 84.33 83.25 83.81 - - 84.25 84.4 82.83
0.4 - - - 83.86 83.81 - - - 83.38 82.76
0.5 - - - - 84.11 - - - - 82.26
Similarly, we evaluate the performance of SGCL-T by varying the number of iteration $K$, while keeping $\tau=0.3$ and $\mu=-0.4$ fixed. The evaluation is performed on both the Cora and Pubmed datasets, with the corresponding accuracy and runtime for different values of $K$. The results are summarized in Table 14.
Table 14: Performance of SGCL-T with fixed $\tau=0.3$ and $\mu=-0.4$ and different numbers of iteration $K$.
K 1 2 3 4 5 6 7
Cora Accuracy 84.25 84.47 83.55 84.37 84.11 84.26 84.11
Time (s) 0.0849 0.1242 0.1727 0.2183 0.2758 0.3211 0.3734
Pubmed Accuracy 84.34 84.25 84.1 83.42 83.35 83.76 83.76
Time (s) 0.6179 1.0517 1.4551 1.9107 2.2914 2.6974 3.1376
From the ablation study, we observe that the highest accuracy across different benchmarks is achieved with varying hyperparameter values. However, for consistency, we set $\tau=0.3$, $\mu=-0.4$, and $K=2$ for all experiments. It’s also worth noting that as $K$ increases, computation time per epoch significantly rises for both datasets, highlighting a trade-off between accuracy and efficiency.
F.3.2 Bilateral smoothing
In the bilateral smoothing approach, the $\sigma_{\text{int}}$ parameter controls how sensitive the bilateral filter is to intensity differences. In this context, it determines how sharply the filter distinguishes between different intensities. When $\sigma_{\text{int}}$ increases, the sensitivity of the filter to differences in intensity decreases, causing it to treat all intensity values more similarly. This makes the bilateral filter act more like a standard Gaussian filter, which smooths uniformly without considering intensity differences. Similarily, the $\sigma_{\text{spa}}$ parameter determines the extent to which spatial proximity affects the bilateral smoothing process. It controls how much weight is given to neighboring nodes based on their distance in the graph. When $\sigma_{\text{spa}}$ is small, only nodes that are very close to each other have a strong influence on each other during smoothing. This preserves fine details and small features. As $\sigma_{\text{spa}}$ increases, the filter begins to smooth over larger distances, meaning that larger features in the graph get smoothed out. In this case, even nodes that are farther apart will start influencing each other more, resulting in a broader, more generalized smoothing effect.
For SGCL-B, we conduct an ablation study on two key hyperparameters: $\sigma_{\text{init}}$ and $\sigma_{\text{spa}}$. The study is carried out on two datasets, Cora and Pubmed, with varying values for both hyperparameters. The results are presented in Table 15, where we evaluate model performance for different combinations of $\sigma_{\text{init}}$ and $\sigma_{\text{spa}}$.
Table 15: Performance of SGCL-B with different combinations of $\sigma_{\text{init}}$ and $\sigma_{\text{spa}}$.
Cora Pubmed
$\sigma_{\text{spa}}$ 0.01 0.05 0.1 0.15 0.2 0.01 0.05 0.1 0.15 0.2
$\sigma_{\text{init}}$ 1 84.48 84.72 84.87 83.96 84.26 83.08 84.11 84.25 83.12 82.14
2 84.37 84.16 84.78 84.16 83.4 83.25 84.23 84.1 82.49 83.96
3 84.47 82.99 84.21 84.21 84.47 84.22 84.26 84.17 83.26 83.28
4 84.47 84.37 83.96 83.91 83.15 83.86 83.92 84.09 83.56 83.87
5 84.26 83.86 85.28 84.01 83.35 83.63 83.89 83.82 83.24 83.8
These results indicate that the choice of hyperparameters impacts the performance of the model, and optimal settings vary across different datasets. Additionally, smaller values of $\sigma_{\text{spa}}$ tend to preserve local graph features better, while larger values induce more smoothing, which affects accuracy differently across datasets.
F.3.3 Diffusion-based smoothing
In this approach, the diffusion rate parameter $\eta$ controls the speed of smoothing. Larger values lead to faster diffusion and more aggressive smoothing, while smaller values retain more of the original structure. The parameter $K$ controls how long the smoothing process continues. More iterations lead to a more globally smooth result, while fewer iterations keep the smoothing more localized. Essentially, diffusion-based smoothing gradually spreads information from each node to its neighbors, helping to equalize values across the graph. The choice of $\eta$ and $K$ allows for fine control over how quickly and broadly this smoothing occurs, making it a flexible approach for graph-based data processing.
In the ablation study for SGCL-D, we examine the impact of two key hyperparameters: $\eta$ (diffusion rate) and $K$ (number of iterations) on the Cora and Pubmed datasets. The results are summarized in Table 16.
Table 16: Performance of SGCL-D with different combinations of $\eta$ and $K$.
Cora Pubmed
$\eta$ 0.01 0.03 0.1 0.2 0.5 0.01 0.03 0.1 0.2 0.5
$K$ 1 84.26 84.67 84.12 84.22 84.11 84.23 84.29 84.17 83.17 83.46
2 84.52 84.47 84.53 84.11 84.23 84.11 85.12 83.65 84.12 83.25
3 84.33 84.17 84.23 83.9 83.87 84.63 85.17 84.27 83.63 84.04
4 84.85 84.52 83.86 83.27 83.66 84.55 84.29 83.66 83.26 83.28
5 84.97 84.28 83.54 83.64 83.78 84.57 83.88 83.46 83.7 83.68
On Cora, the highest accuracy (84.97(84.97( 84.97%) is achieved with $\eta=0.01$ and $K=5$, indicating that slower diffusion and more iterations tend to yield better results. For Pubmed, the best accuracy ($85.17$%) is achieved with $\eta=0.03$ and $K=3$, showing that moderate diffusion rates paired with a balanced number of iterations deliver optimal results. These results highlight a trade-off between diffusion speed and the number of iterations: smaller $\eta$ values combined with more iterations generally retain more local details, while larger $\eta$ values spread smoothing effects faster but may reduce accuracy due to over-smoothing.
Appendix G Limitations and future directions
As discussed, contrasting with true hard negatives—nodes with different labels from the anchor but located nearby—can improve model performance by providing strong gradient signals and enhancing contrast. However, in many benchmarks, particularly in homophilic graphs, these nearby nodes are likely to share the same label as the anchor, making them potential false negatives. This issue can undesirably push away semantically similar samples, resulting in performance degradation. The proposed SGCL methods effectively mitigate the harmful impact of false negative nodes. However, this also inevitably reduces the influence of potential true hard negative nodes. Despite this trade-off, the overall reduction in false negatives is expected to have a more significant positive effect than the reduced influence of true hard negatives, as supported by our analytical analysis.
As the first to introduce smoothed positive/negative pairs for graph contrastive learning, we explored the development of a stable and effective learnable smoothing objective. However, we found that a straightforward learnable solution is challenging to train since a learnable smoothing objective can make the overall loss unstable. Consequently, the basic version did not yield performance improvements. Nonetheless, we recognize this as a crucial direction for future enhancement and view it as an exciting avenue for further research.